OpenStudy (anonymous):

what is the equation of the line in standard form that point (1,5) and is parallel to the line y=-3x-3

OpenStudy (anonymous):

y-5=-3(x-1) y-5=-3x+3 y=-3x+8

OpenStudy (anonymous):

a. x+3y=-6 b. x-3y=0 c. 3x+y=6 d. 3x+y=8 Is it d ?